OpenStudy (anonymous):

HELP ! - y/3 (>=) -5 A) y (<=) 15 B) y (<=) -8 C) y (>=) -15 D) y (>=) 15

Nnesha (nnesha):

this is same like last one

Nnesha (nnesha):

you just have to multiply 3 both side

sammixboo (sammixboo):

Correct. \(\rm\dfrac{y}{3}\ge -5\) \(\rm\dfrac{y}{3}\color{blueviolet}{\times3}\ge -5\color{blueviolet}{\times 3}\) \(\rm y\ge -5\color{blueviolet}{\times 3}\) What is -5 x 3 REMEMBER - When multiplying by negative (same with dividing) you always flip the inequality sign! So we have this \(\rm y\le -5\color{blueviolet}{\times 3}\)

Nnesha (nnesha):

what you get ??

sammixboo (sammixboo):

Oh and... \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\) Once again =)

OpenStudy (anonymous):

thx !

sammixboo (sammixboo):

So, what did you get?

OpenStudy (anonymous):

-15 ?

sammixboo (sammixboo):

Correct, \(\rm y \le -15\)

OpenStudy (anonymous):

thx ! :D